# Characterizing the Post-Red Supergiant binary system AFGL 4106 and its complex nebula with SPHERE/VLT

### v1

@@ Abstract
Context. Mass loss in evolved massive stars plays a critical role in shaping their circumstellar environments and enriching the interstellar medium. In binary systems, stellar interactions can further complicate this process, affecting both stellar evolution, stellar yields and nebular morphology.
Aims. We aim to characterize the physical and morphological properties of the binary system AFGL 4106, composed of two evolved massive stars. Understanding its mass-loss processes and circumstellar environment offers insight into the late stages of stellar evolution in massive binary systems.
Methods. We obtained high-angular–resolution, high-contrast imaging using VLT/SPHERE with ZIMPOL (optical) and IRDIS (near-infrared) across multiple filters. We used aperture photometry to extract the spectral energy distributions (SEDs) of each star, and applied radiative transfer modelling to study the system and its surrounding dusty environment.
Results. The observations resolve both components of the binary and unveil a complex, dusty nebula featuring asymmetric structures and cavities. SED fitting yields stellar temperatures of T1 = 6723 196 K and T2 = 3394 264 K, along with bolometric luminosities of L L⊙ and L L⊙. These values support the classification of the primary as being in a post-red supergiant (post-RSG) phase and the secondary as an active red supergiant (RSG). The luminosity ratio, combined with the inferred radii, indicates that both stars are at close yet distinct stages of their evolution. The binary is surrounded by an extended shell whose asymmetric morphology and large-scale features suggest interaction with the stellar winds and interstellar medium (ISM), and possibly the presence of a third, undetected companion.
Conclusions. These observations provide the first resolved view of AFGL 4106’s system and its dusty envelope. Our analysis sets constraints on the physical properties and evolutionary status of the system. This work contributes to understanding mass-loss processes in massive binaries and the shaping of nebulae around evolved stars.
@@
Mass loss from massive stars governs the chemical enrichment of the interstellar medium. Post-Red Supergiants (post-RSGs) are prime laboratories to study the morphology, composition and kinematics of recently ejected material, and to reconstruct mass-loss histories, since they can undergo very strong outflows (up to ; Mauron & Josselin (2011)). The class is small: only a few well-studied examples exist, such as IRC +10420 (Koumpia et al., 2022) and the ”Fried Egg” nebula (Lagadec et al., 2011; Koumpia et al., 2020). For a recent overview of post-RSGs and their circumstellar environments see Gordon & Humphreys (2019), and for RSGs, their mass loss and subsequent evolution see van Loon (2025).
@@
The complex nebula shapes observed around such objects can be due to several factors such as magnetic fields or rotation, but it is often thought that the presence of companion(s) is the main source of the angular momentum excess necessary for their shaping (Balick & Frank, 2002). Observational studies and models have shown that a large proportion of stars, especially solar-type and massive stars, reside in binary or multiple systems (Whitworth & Lomax, 2015; Fuhrmann et al., 2017; Merle, 2024). Companions must have an impact on the mass loss compositions and rate from massive stars. Studying the mass loss history of Post-Red Supergiants and the influence of binaries is a key to understand matter ejection from massive stars before the supernova explosion.
@@
AFGL 4106 provides an excellent case study for investigating how binarity influences mass loss in massive stars. It was examined in the late 1990s by Molster et al. (1999) and van Loon et al. (1999), both of which reached similar conclusions: the system is binary, consisting of an evolved star—likely a post-Red Supergiant—and a less evolved companion, embedded within a circumbinary nebula produced by one or both stars. The primary (star 1) is estimated to have a temperature of 7250 250 K, while the companion (star 2) has a temperature of 3750 250 K . The binary system was first resolved in 2006 using Hubble Space Telescope (HST) observations, leading to a separation measurement of 0.3” with a position angle P.A. = 270° (Bobrowsky et al., 2006).
@@
The Spectro-Polarimetric High-contrast Exoplanet REsearch (SPHERE) on the Very Large Telescope provides an exquisite 20 milliarcsec angular resolution in the optical and 50 milliarcsec in the near-infrared, enabling to better characterize the central stars of the system. SPHERE’s coronagraph and polarimetric capabilities enable a deep and detailed study of the nebula around the central binary system, allowing, for the first time, to study the mass-loss history of a resolved binary system including a Post-Red supergiant star.
@@
In this study, we present a detailed analysis of the evolved massive binary system AFGL 4106, based on high-angular-resolution, high-contrast observations obtained with VLT/SPHERE in both direct and polarimetric imaging modes across multiple optical and near-infrared wavelengths. These data reveal the morphology and scattering properties of the surrounding dusty nebula, bringing constraints on the mass-loss history of the system.
@@
This paper is structured as follows: in Section 2, we describe the SPHERE observations and data reduction. Section 3 presents the morphological and photometric analysis, including the spectral energy distributions of each component and discuss the radiative transfer modelling used to constrain the stellar and dust parameters. In Section 4, we interpret the results in the context of stellar evolution, mass-loss history, and nebular shaping mechanisms. Finally, conclusions are summarized in Section 5.
2 Observation and Data reductions
@@ 2.1 SPHERE observations
Observations were carried out with the extreme adaptive optics imager and polarimeter SPHERE (Beuzit et al., 2019) installed at the Unit Telescope 3 Nasmyth focus of the Very Large Telescope (VLT). Its adaptive optics module SAXO (Fusco et al., 2006) provides a correction from the atmosphere turbulence to reach the highest angular resolution possible. The main component of the adaptive optics system is the Shack-Hartmann wave front sensor (WFS), with 4141 actuators working at a frequency of 1380 Hz on the deformable mirror.
We obtained intensity images of AFGL 4106 using the SPHERE instruments ZIMPOL (optical, 0.5-0.9 m) and IRDIS (near-infrared, 0.95-2.32 m). To enhance contrast in the ZIMPOL data, we acquired polarimetric maps, as dust-scattered light is polarized while stellar emission is not, revealing the circumstellar environment at flux levels several orders of magnitude fainter than the stars. For selected optical filters, we combined polarimetry with a 77.5 mas coronagraph to further suppress the primary star’s intensity and improve contrast. Additional data were obtained without a coronagraph to probe the innermost regions of the system. IRDIS observations were performed in classical imaging mode, since no coronagraph was large enough to mask both stars; we increased the exposure time and allowed the central region to saturate in some IRDIS frames to better detect the nebula. A summary of the SPHERE observations is provided in Tab. 3.
Both ZIMPOL and IRDIS data were reduced using the SPHERE data reduction center (Delorme et al., 2017) which makes use of the Zurich ZIMPOL pipeline (Schmid et al., 2018) and the SPHERE pipeline (Pavlov et al., 2008) for the basic reduction steps. The SPHERE data center makes use of custom designed python routines to provide the ZIMPOL polarisation map. In polarization mode, ZIMPOL measure the Stokes parameter Q and U, the polarization maps are then computed using the following equations:
and where is the polarized intensity, the total intensity, the polarization angle, and the normalized Stokes parameters, and the total intensity measured in the Q and U cycles respectively, and , , and are the Stokes parameters measured in each half cycle.
@@ 2.2 Auxiliary photometry
The spectral energy distribution (SED) of the system is affected by the properties of the two stars and the interaction of their light with dust around them. To construct the SED of AFGL 4106, we retrieved archival multi-wavelength photometric data covering both the optical and infrared ranges. This broad spectral coverage allows us to characterize the contributions from both the stellar components and the surrounding dust. Additionally, we incorporated ISO-SWS spectra (Sloan et al., 2003) to complement the infrared portion of the SED. The spectra were integrated over ISO filter response curves to extract corresponding photometric points. The data we used are presented in Table 4.
@@ 3.1 The central stars
From Hubble Space Telescope (HST) imaging, the separation between the two central stars has been estimated to 0.3” (Bobrowsky et al., 2006) with a position angle of 270∘ for the secondary. However, the theoretical diffraction limit () of HST at 1.644 m is 0.141”. This implies that the separation of 0.3” is just above the Nyquist limit. With our SPHERE ZIMPOL images, we reach a theoretical diffraction limit of 0.014” in V band (Fig. 1), which is about one order of magnitude better than HST resolution. This allows us to get a better estimation of the separation. We performed a photocenter fit on each components in the V band image in order to measure the separation. For this, we used the Iteratively Weighted Centre of Gravity (IWCoG) technique (Vyas et al., 2009), where the photocenter is defined as:
Where (x, y) are the photocenter coordinates, X is the position of the pixel, I the value associated with the pixel and W a Gaussian weight, calculated using a radial profile of the star. We finally find a separation of 0.272 0.017 arcsec with a position angle P.A. 90°. At a distance of 3.19 kpc (see Sec. 4.1), this corresponds to a physical separation of 865 102 au. Our measurement differs from that reported by Bobrowsky et al. (2006). This discrepancy arises from the choice of the reference star, which introduces a 180° ambiguity in the position angle. They indeed used star 2 as a reference, while we used star 1 (the one in the east in all images).
@@ 3.2 The Nebula
3.2.1 ZIMPOL optical imaging polarimetry
With its narrow field of view () and high angular resolution (about 20 mas, Schmid et al. (2018)), observations with the ZIMPOL instrument reveal the morphology of the inner part of the nebula. To enhance the contrast, we used a coronagraph (seen as a black circle with spiders) and polarimetric measurements. The light from the star is unpolarized, while dust scatters this light. Images in polarized light enable to partially mask the central star and reveal the dust around it. Figure 3 presents the polarized intensity image obtained with the ZIMPOL I_PRIM filter (central wavelength 789.7 nm).
The central circular shape, with a diameter of 1 arcsec, is the adaptive optics correction halo, and therefore, not physical. The ZIMPOL polarized intensity image (Fig. 3) highlights two prominent cavities, referred to as Holes 1 and 2. These cavities, each measuring approximately in length and in width, are positioned symmetrically on either side of a bright North East-South West axis.
In the polarized intensity images, the two stars appear embedded in a non-uniform cavity of 2 arcsec in diameter, including the two voids described before. The inner edges of this cavity are clearly nonspherical and a dichotomy is observed between the North-West rim where four loops of diameter 0.3 arcsec are observed and the South-East displays 5 cometary-like features pointing toward the center.
The external parts of the observed nebula are more uniform, albeit still clumpy.
@@ 3.2.2 IRDIS near-infrared imaging
The IRDIS intensity images reveal the outer parts of the nebula. As no coronagraph were large enough to mask the central stars, the central stars were heavily saturated in some sequences to reveal the nebula.
The nebula has a diameter of approximately . It exhibits a variety of intricate structures at multiple scales. Inside the nebula, the saturated stars appear has black dots. The two double lines crossing the stars are the spider of M2. The 1.5 arcsec diameter circle is the AO correction halo. Two filaments are observed on the South-East and West sides of the nebula, about 2 arcsec from the center, clearly visible in (Fig. 4). Another one is present in the North-East, about 3 arcsec away from the center. Finally one is present in the North-West about 2 arcsec away from the center. The gaps between the filaments are not empty, but rather clumpy.
At larger scale, on the IRDIS images, we observe sharp triangle-shaped lobes coming out radially from the central star. They are observed at various distances between 1.5 and 2.0 arcsec from the central star. They appear to be sharper towards the south and west of the nebula.
They are similar to the lobes observed on CPD-568032 (Chesneau et al., 2006)or in the starfish nebula (Sahai et al., 2005) and are also similar to the sporadic, clumpy ejecta imaged around VY CMa (Humphreys et al., 2021). Mostly visible in filter J (Fig. 4) and Continuum J (Fig. 18), these lobes are not visible at larger wavelengths. The most prominent structure is the southern one denoted as S Lobe on Fig. 2, which is 2” long. The other lobes are smaller and all 0.5” long.
@@
The iso-intensity map of the nebula displays variations in the shape of the ejecta (Fig. 5(a)). Looking up at the eccentricity of these iso-intensity curve, where a is the semi-major axis and b the semi minor axis, seems to change with distance to the central star. We observe oscillations of pseudo-period 2.560 arcsec, given that we only have 1.5 periods. The gap observed between 1.8 and 2.2 arcseconds corresponds to a sharp variation of ellipticity, likely caused by a rotation of the elliptical contours by , resulting in a swap between the semi-major and semi-minor axes. This suggest a change in the ejecta preferential direction across time
Assuming a typical stellar wind velocity of 20-30 km/s (Molster et al., 1999) and a distance of 3.19 kpc (see discussion in Section 4.1), we estimate oscillation periods ranging from approximately 1300 to 1950 years, depending on the adopted wind speed.
@@ 3.3 Spectral energy distribution of both stars
Using the first resolved observations of both components, we performed aperture photometry on our SPHERE data to isolate the SED of each star in the system across various optical and near-infrared wavelengths. Flux ratios were derived from the SPHERE data by integrating the flux within circular apertures of radius , corresponding to the diffraction-limited core and ensuring that the full dispersed flux is captured. Among the 16 available filters, only 5 provide a suitable combination of PSF quality and unsaturated stellar images, allowing us to perform aperture photometry on both stars.
Since SPHERE data are not flux-calibrated, the total flux for each star was obtained from photometric auxiliary data. These fluxes were corrected from the interstellar reddening using PyAstronomy unred routine (Czesla et al., 2019) making use of the Fitzpatrick (1999) parametrization to estimate the wavelength dependence of interstellar extinction. This parametrization uses spectral observation to derive the shape of the extinction curve, fitted to the following formula:
where E(-V) is the extinction at wavelength , E(B-V) the extinction in the B and V bands and :
@@
It is worth noting that that GSP-Spec has analysed a Gaia/RVS spectrum that is actually a combined spectrum of the unresolved binary system, which may slightly affect the atmospheric parameter determination. Moreover, as AFGL 4106 is likely variable, the derived extinction is estimated from an average spectrum obtained over the 2.5 years of Gaia/DR3 observations. Using our adopted stellar temperature ( K, see (Sec. 3.4)), the expected intrinsic colour becomes mag, leading to a slightly higher reddening, . This difference remains small and does not significantly affect the other derived parameters, such as luminosity, but should be mentioned.
To facilitate comparison with the extinction from Molster et al. (1999), we converted to using the extinction coefficients and from the ESA/Gaia photometric system documentation (Riello et al., 2021). Therefore, we have:
Adopting the standard value of for the Milky Way interstellar medium, the total extinction reported by Gaia corresponds to mag. This value is consistent with that reported by Molster et al. (1999), considering that they found an extinction due to the circumstellar material of mag. This leads to a total extinction in agreement with the one used by Molster et al. (1999), and in the following analysis, we adopt the same interstellar reddening: E(B-V) = 1.07 mag.
| SPHERE Filter | (m) | 111 erg/s/cm2 | |
@@
For the first time, we resolved the binary system in the core of AFGL 4106 at multiple wavelengths in the optical and near-infrared, allowing us to build SEDs for both objects. The five photometric data points clearly indicate that the primary star is hotter and more luminous than its companion.
To have a better estimation of the stellar properties, we performed radiative transfer modelling of the system and its surrounding dust shell, as described in the next section.
3.4 Stars properties
Fig. 6 give us an overview of the SED of each component of the system in 5 bands. We used the 1D radiative transfer code DUSTY (Version 4, Ivezic & Elitzur (1997)) to model the SED of the system, adopting the dust composition identified by Molster et al. (1999). The model aims to fit the photometric data obtained from VizieR 222https://vizier.cds.unistra.fr, as well as photometric points constructed by convolving the ISO-SWS spectrum with the ISO filter profiles (Tab. 4) . We use the standard MRN dust size distribution (Mathis et al., 1977) and we opted for the solution for radiatively driven winds. The best fit is determined with a minimization of the using Monte Carlo Markov Chain. We created synthetic fluxes by convolving the model spectrum with the filter profiles of each observational instrument. We then used the defined by (Yang et al., 2023):
where () = F()/F(K) is the flux at a specific wavelength normalized to the flux in the K band, N is the number of photometric data points (5 in the case of Tab. 1 and 47 in the case of Tab. 4), and p is the number of free parameters (5 to 7). This modified balances the contributions from both optical and infrared fluxes by effectively flattening the spectrum. In contrast, a standard tends to underweight the infrared data due to the steep decline in flux at longer wavelengths of the Planck function.
We initially used the SEDs of each component, as shown in Fig. 6, to estimate the stellar temperatures through a first MCMC run. Based on these results, we then constrained the temperatures to narrow intervals centered on the initial estimates. Using the full photometric dataset available from VizieR, we subsequently refined the model to determine the properties of the circumstellar shell and the relative luminosities of the two components in a second MCMC run.111To facilitate the reusability of our method, we developed DustyPY https://github.com/gtomass/DustyPY.git, a Python library available on GitHub.

### v2

# Characterising the post-red supergiant binary system AFGL 4106 and its complex nebula with SPHERE/VLT

@@ Abstract
Context. Mass loss in evolved massive stars plays a critical role in shaping their circumstellar environments and enriching the interstellar medium. In binary systems, stellar interactions can further complicate this process, affecting stellar evolution, stellar yields, and nebular morphology.
Aims. We aim to characterise the physical and morphological properties of the binary system AFGL 4106, which is composed of two evolved massive stars. Understanding its mass-loss processes and circumstellar environment offers insight into the late stages of stellar evolution in massive binary systems.
Methods. We obtained high-angular-resolution, high-contrast imaging using VLT/SPHERE with ZIMPOL (optical) and IRDIS (near-infrared) across multiple filters. We used aperture photometry to extract the spectral energy distributions (SEDs) of each star and applied radiative-transfer modelling to study the system and its surrounding dusty environment.
Results. The observations resolve both components of the binary and unveil a complex, dusty nebula featuring asymmetric structures and cavities. SED fitting yields stellar temperatures of T1 = 6723 196 K and T2 = 3394 264 K, along with bolometric luminosities of L L⊙ and L L⊙. These values support the classification of the primary as being in a post-red-supergiant (post-RSG) phase and the secondary as an active red supergiant (RSG). The luminosity ratio, combined with the inferred radii, indicates that both stars are at close yet distinct stages of their evolution. The binary is surrounded by an extended shell whose asymmetric morphology and large-scale features suggest interaction with the stellar winds and interstellar medium (ISM), and possibly the presence of a third, undetected companion.
Conclusions. These observations provide the first resolved view of AFGL 4106’s system and its dusty envelope. Our analysis sets constraints on the physical properties and evolutionary status of the system. This work contributes to our understanding of mass-loss processes in massive binaries and the shaping of nebulae around evolved stars.
@@
Mass loss from massive stars governs the chemical enrichment of the interstellar medium. Post-red supergiants (post-RSGs) are prime laboratories in which to study the morphology, composition, and kinematics of recently ejected material, and to reconstruct mass-loss histories, since they can undergo very strong outflows (up to ; Mauron & Josselin (2011)). The class is small: only a few well-studied examples exist, such as IRC +10420 (Koumpia et al., 2022) and the ’Fried Egg’ nebula (Lagadec et al., 2011; Koumpia et al., 2020). For a recent overview of post-RSGs and their circumstellar environments, see Gordon & Humphreys (2019); and for RSGs, their mass loss, and subsequent evolution, see van Loon (2025).
@@
The complex nebula shapes observed around such objects can be caused by several factors, such as magnetic fields or rotation, but it is often thought that the presence of one or more companions is the main source of the angular momentum excess necessary for their shaping (Balick & Frank, 2002). Observational studies and models have shown that a large proportion of stars –especially solar-type and massive stars– reside in binary or multiple systems (Whitworth & Lomax, 2015; Fuhrmann et al., 2017; Merle, 2024). Companions must have an impact on the mass-loss compositions and rate from massive stars. Studying the mass loss history of post-RSGs and the influence of binaries is key to understanding matter ejection from massive stars before the supernova explosion.
@@
AFGL 4106 provides an excellent case study for investigating how binarity influences mass loss in massive stars. It was examined in the late 1990s by Molster et al. (1999) and van Loon et al. (1999), both of which reached similar conclusions: the system is a binary, consisting of an evolved star –likely a post-RSG– and a less evolved companion, embedded within a circumbinary nebula produced by one or both stars. The primary (star 1) is estimated to have a temperature of 7250 250 K, while the companion (star 2) has a temperature of 3750 250 K. The binary system was first resolved in 2006 using Hubble Space Telescope (HST) observations, leading to a separation measurement of 0.3” with a position angle (P.A.) of 270° (Bobrowsky et al., 2006).
@@
The Spectro-Polarimetric High-contrast Exoplanet REsearch (SPHERE) on the Very Large Telescope provides an exquisite 20 milliarcsec angular resolution in the optical and 50 milliarcsec in the near-infrared, enabling us to better characterise the central stars of the system. SPHERE’s coronagraph and polarimetric capabilities enable a deep and detailed study of the nebula around the central binary system, allowing, for the first time, a study of the mass-loss history of a resolved binary system including a post-RSG star.
@@
In this paper, we present a detailed analysis of the evolved massive binary system AFGL 4106; it is based on high-angular-resolution, high-contrast observations obtained with VLT/SPHERE in both direct and polarimetric imaging modes across multiple optical and near-infrared wavelengths. These data reveal the morphology and scattering properties of the surrounding dusty nebula, putting constraints on the mass-loss history of the system.
@@
This paper is structured as follows. In Section 2, we describe the SPHERE observations and data reduction. Section 3 presents the morphological and photometric analysis, including the spectral energy distributions of each component; and in it we discuss the radiative-transfer modelling used to constrain the stellar and dust parameters. In Section 4, we interpret the results in the context of stellar evolution, mass-loss history, and nebular shaping mechanisms. Finally, our conclusions are summarised in Section 5.
2 Observations and data reductions
@@ 2.1 SPHERE observations
Observations were carried out with the extreme adaptive optics imager and polarimeter SPHERE (Beuzit et al., 2019) installed at the Unit Telescope 3 Nasmyth focus of the Very Large Telescope (VLT). Its adaptive optics module SAXO (Fusco et al., 2006) provides a correction from the atmosphere turbulence to reach the highest angular resolution possible. The main component of the adaptive optics system is the Shack-Hartmann wave-front sensor (WFS), with 4141 actuators working at a frequency of 1380 Hz on the deformable mirror.
We obtained intensity images of AFGL 4106 using the SPHERE instruments ZIMPOL (optical, 0.5-0.9 m) and IRDIS (near-infrared, 0.95-2.32 m). To enhance contrast in the ZIMPOL data, we acquired polarimetric maps, as dust-scattered light is polarised, while stellar emission is not, revealing the circumstellar environment at flux levels several orders of magnitude fainter than the stars. For selected optical filters, we combined polarimetry with a 77.5 mas coronagraph to further suppress the primary star’s intensity and improve contrast. Additional data were obtained without a coronagraph to probe the innermost regions of the system. IRDIS observations were performed in classical imaging mode, since no coronagraph was large enough to mask both stars; we increased the exposure time and allowed the central region to saturate in some IRDIS frames to better detect the nebula. A summary of the SPHERE observations is provided in Tab. 3.
Both ZIMPOL and IRDIS data were reduced using the SPHERE data-reduction centre (Delorme et al., 2017), which makes use of the Zurich ZIMPOL pipeline (Schmid et al., 2018) and the SPHERE pipeline (Pavlov et al., 2008) for the basic reduction steps. The SPHERE data centre makes use of custom-designed python routines to provide the ZIMPOL polarisation map. In polarisation mode, ZIMPOL measures the Stokes parameters Q and U; the polarisation maps are then computed using the following equations:
and where is the polarized intensity, the total intensity, the polarization angle, and the normalized Stokes parameters, and the total intensity measured in the Q and U cycles, respectively, and , , and the Stokes parameters measured in each half cycle.
@@ 2.2 Auxiliary photometry
The spectral energy distribution (SED) of the system is affected by the properties of the two stars and the interaction of their light with dust around them. To construct the SED of AFGL 4106, we retrieved archival multi-wavelength photometric data covering both the optical and infrared ranges. This broad spectral coverage allows us to characterise the contributions from both the stellar components and the surrounding dust. Additionally, we incorporated ISO-SWS spectra (Sloan et al., 2003) to complement the infrared portion of the SED. The spectra were integrated over ISO filter response curves to extract corresponding photometric points. The data we used are presented in Table 4.
@@ 3.1 The central stars
From Hubble Space Telescope (HST) imaging, the separation between the two central stars has been estimated as 0.3” (Bobrowsky et al., 2006) with a position angle of 270∘ for the secondary. However, the theoretical diffraction limit () of HST at 1.644 m is 0.141”. This implies that the separation of 0.3” is just above the Nyquist limit. With our SPHERE ZIMPOL images, we reached a theoretical diffraction limit of 0.014” in the V band (Fig. 1), which is about one order of magnitude better than HST resolution. This allows us to obtain a better estimation of the separation. We performed a photocentre fit on each components in the V-band image in order to measure the separation. For this, we used the iteratively-weighted centre-of-gravity (IWCoG) technique (Vyas et al., 2009), where the photocentre is defined as
where (x, y) are the photocentre coordinates, X is the position of the pixel, I the value associated with the pixel, and W a Gaussian weight, calculated using a radial profile of the star. We finally find a separation of 0.272 0.017 arcsec with a position angle of 90°. At a distance of 3.19 kpc (see Sect. 4.1), this corresponds to a physical separation of 865 102 au. Our measurement differs from that reported by Bobrowsky et al. (2006). This discrepancy arises from the choice of the reference star, which introduces a 180 degree° ambiguity in the position angle. They indeed used star 2 as a reference, while we used star 1 (the one in the east in all images).
@@ 3.2 The nebula
3.2.1 ZIMPOL optical-imaging polarimetry
With its narrow field of view () and high angular resolution (about 20 mas, Schmid et al. (2018)), observations with the ZIMPOL instrument reveal the morphology of the inner part of the nebula. To enhance the contrast, we used a coronagraph (seen as a black circle with spiders) and polarimetric measurements. The light from the star is unpolarised, while dust scatters this light. Images in polarised light enabled us to partially mask the central star and reveal the dust around it. Figure 3 presents the polarised-intensity image obtained with the ZIMPOL I_PRIM filter (central wavelength: 789.7 nm).
The central circular shape, with a diameter of 1 arcsec, is the adaptive optics correction halo and, therefore, not physical. The ZIMPOL polarised-intensity image (Fig. 3) highlights two prominent cavities, which are referred to as holes 1 and 2. These cavities, each measuring approximately in length and in width, are positioned symmetrically on either side of a bright northeast-southwest axis.
In the polarised-intensity images, the two stars appear embedded in a non-uniform cavity of 2 arcsec in diameter, including the two voids described before. The inner edges of this cavity are clearly non-spherical, and a dichotomy is observed between the northwest rim, where four loops of 0.3 arcsec in diameter are observed, and the southeast displays five comet-like features pointing towards the centre. The external parts of the observed nebula are more uniform, albeit still clumpy.
@@ 3.2.2 IRDIS near-infrared imaging
The IRDIS intensity images reveal the outer parts of the nebula. As no coronagraphs were large enough to mask the central stars, the central stars were heavily saturated in some sequences to reveal the nebula.
The nebula has a diameter of approximately . It exhibits a variety of intricate structures at multiple scales. Inside the nebula, the saturated stars appear has black dots. The two double lines crossing the stars represent the spider of M2. The 1.5 arcsec diameter circle is the AO correction halo. Two filaments are observed on the southeast and west sides of the nebula, about 2 arcsec from the centre, clearly visible in (Fig. 4). Another one is present in the northeast, about 3 arcsec away from the centre. Finally one is present in the northwest about 2 arcsec away from the centre. The gaps between the filaments are not empty, but rather clumpy.
On a larger scale, in the IRDIS images we observe sharp, triangle-shaped lobes exiting the central star radially. They are observed at various distances between 1.5 and 2.0 arcsec from the central star. They appear to be sharper towards the south and west of the nebula.
They are similar to the lobes observed on CPD-568032 (Chesneau et al., 2006) or in the starfish nebula (Sahai et al., 2005) and are also similar to the sporadic, clumpy ejecta imaged around VY CMa (Humphreys et al., 2021). Mostly visible in filter J (Fig. 4) and continuum J (Fig. 18), these lobes are not visible at larger wavelengths. The most prominent structure is the southern one denoted as S lobe on Fig. 2, which is 2” long. The other lobes are smaller, and all are 0.5” long.
@@
The iso-intensity map of the nebula displays variations in the shape of the ejecta (Fig. 5(a)). Looking up at the eccentricity of these iso-intensity curve, , where a is the semi-major axis and b the semi-minor axis, and seems to change with distance to the central star. We observe oscillations with a pseudo-period of 2.560 arcsec, given that we only have 1.5 periods. The gap observed between 1.8 and 2.2 arcseconds corresponds to a sharp variation of ellipticity, likely caused by a rotation of the elliptical contours by , resulting in a swap between the semi-major and semi-minor axes. This suggest a change in the ejecta preferential direction across time.
Assuming a typical stellar-wind velocity of 20-30 km/s (Molster et al., 1999) and a distance of 3.19 kpc (see discussion in Sect. 4.1), we estimate oscillation periods ranging from approximately 1300 to 1950 years, depending on the adopted wind speed.
@@ 3.3 Spectral energy distribution of both stars
Using the first resolved observations of both components, we performed aperture photometry on our SPHERE data to isolate the SED of each star in the system across various optical and near-infrared wavelengths. Flux ratios were derived from the SPHERE data by integrating the flux within circular apertures of radius , corresponding to the diffraction-limited core and ensuring that the full dispersed flux is captured. Among the 16 available filters, only five provide a suitable combination of PSF quality and unsaturated stellar images, allowing us to perform aperture photometry on both stars.
Since SPHERE data are not flux-calibrated, the total flux for each star was obtained from photometric auxiliary data. These fluxes were corrected from the interstellar reddening using the PyAstronomy unred routine (Czesla et al., 2019), making use of the Fitzpatrick (1999) parametrisation to estimate the wavelength dependence of interstellar extinction. This parametrisation uses spectral observation to derive the shape of the extinction curve, fitted to the following formula:
where E(-V) is the extinction at wavelength , E(B-V) the extinction in the B and V bands, and
@@
It is worth noting that that GSP-Spec has analysed a Gaia/RVS spectrum that is actually a combined spectrum of the unresolved binary system, which may slightly affect the atmospheric parameter determination. Moreover, as AFGL 4106 is likely variable, the derived extinction is estimated from an average spectrum obtained over the 2.5 years of Gaia/DR3 observations. Using our adopted stellar temperature ( K, see (Sect. 3.4)), the expected intrinsic colour becomes mag, leading to a slightly higher reddening, . This difference remains small and does not significantly affect the other derived parameters, such as luminosity, but should be mentioned.
To facilitate comparison with the extinction from Molster et al. (1999), we converted to using the extinction coefficients and from the ESA/Gaia photometric system documentation (Riello et al., 2021). Therefore, we have
Adopting the standard value of for the Milky Way interstellar medium, the total extinction reported by Gaia corresponds to mag. This value is consistent with that reported by Molster et al. (1999), considering that they found an extinction due to the circumstellar material of mag. This leads to a total extinction in agreement with the one used by Molster et al. (1999). In the following analysis, we adopt the same interstellar reddening: E(B-V) = 1.07 mag. The measurements of the photometry are presented in Tab. 1 and plotted in Fig. 6.
| SPHERE Filter | (m) | 1111 erg/s/cm2 | |
@@
For the first time, we resolved the binary system in the core of AFGL 4106 at multiple wavelengths in the optical and near-infrared, allowing us to build SEDs for both objects. The five photometric data points clearly indicate that the primary star is hotter and more luminous than its companion. To have a better estimation of the stellar properties, we performed radiative-transfer modelling of the system and its surrounding dust shell, as described in the next section.
3.4 Star properties
Figure 6 give us an overview of the SED of each component of the system in five bands. We used the 1D radiative transfer code DUSTY (Version 4, Ivezic & Elitzur (1997)) to model the SED of the system, adopting the dust composition identified by Molster et al. (1999). The model aims to fit the photometric data obtained from VizieR,222https://vizier.cds.unistra.fr as well as photometric points constructed by convolving the ISO-SWS spectrum with the ISO filter profiles (Tab. 4). We used the standard MRN dust size distribution (Mathis et al., 1977) and opted for the solution for radiatively driven winds. The best fit is determined with a minimisation of the using a Markov chain Monte Carlo method. We created synthetic fluxes by convolving the model spectrum with the filter profiles of each observational instrument. We then used the defined by Yang et al. (2023):
where () = F()/F(K) is the flux at a specific wavelength, normalised to the flux in the K band; N is the number of photometric data points (5 in the case of Tab. 1 and 47 in the case of Tab. 4); and p is the number of free parameters (5 to 7). This modified balances the contributions from both optical and infrared fluxes by effectively flattening the spectrum. In contrast, a standard tends to underweight the infrared data due to the steep decline in flux at longer wavelengths of the Planck function.
We initially used the SEDs of each component, as shown in Fig. 6, to estimate the stellar temperatures through a first MCMC run. Based on these results, we then constrained the temperatures to narrow intervals centred on the initial estimates. Using the full photometric dataset available from VizieR, we subsequently refined the model to determine the properties of the circumstellar shell and the relative luminosities of the two components in a second MCMC run.222To facilitate the reusability of our method, we developed DustyPY https://github.com/gtomass/DustyPY.git, a Python library available on GitHub. The final results are presented in Fig. 7 and the model parameters in Table. 3.